# Learning Dynamics of Meta-Learning in Small Model Pretraining

### v1

@@ Abstract
Large language models are powerful but costly. We ask whether meta-learning can make the pretraining of small language models not only better but also more interpretable. We integrate first–order MAML with subset-masked LM pretraining, producing four LLama-style decoder-only models (11M-570M params), and evaluate it on a fundamental NLP task with many settings and real-world applications. Compared with vanilla training, our model (i) reaches the same loss up to 1.6× sooner, (ii) improves F1 on multilingual Universal NER under equal compute, and (iii) makes the training dynamics easy to read: first the network’s representations fan out (“diversify”) and later they collapse into a smaller, shared subspace (“compress”). This two-stage shift shows up as a rise-and-fall in both effective-rank curves and attention-head entropy. The same curves pinpoint which layers specialise earliest and which later reconverge, giving a compact, interpretable signature of meta-adaptation. Code, checkpoints and WandB logs are released.
| pico-maml (Apache 2.0) | |
| pico-maml (Apache 2.0) |
@@
David Demitri Africa††thanks: Corresponding author: dda28@cam.ac.uk Yuval Weiss Paula Buttery Richard Diehl Martinez University of Cambridge
@@ 1 Introduction
Small language models (SLMs) are attractive for privacy and energy reasons, but trail large models partly because they converge slowly and plateau early (Godey et al., 2024; Biderman et al., 2023; Diehl Martinez et al., 2024). As opposed to the common method of brute-force scaling, we explore a different axis: learning rules. First-order Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) promises a learn-to-learn initialization, yet has rarely been applied to decoder models, and its effect on learning dynamics are poorly understood.
We address this by adding meta-learning in model pretraining111Using a lightweight modification of pico-train (Diehl Martinez, 2025), a language model pretraining framework., interleaving ordinary next-token loss (keeps fluency) with 32-way subset-mask (Bansal et al., 2020; Li and Zhang, 2021) episodes (forces rapid binding). Only a tiny MLP head is adapted in the inner loop, so we can track backbone weights without gradient noise. Our contributions are:
@@
Four open SLMs (11M → 570M) trained with this hybrid rule.
@@
A public trainer that logs per-checkpoint singular-value spectra, head entropies and query accuracy to make learning dynamics inspectable.
@@
Evidence that an early "diversify-then-compress" in effective rank predicts final NER F1 improvements.
@@ 2 Related Work
Meta-learning for NLP. (MAML; Finn et al., 2017) is an optimisation-based form of meta-learning that learns an initialisation from which a few gradient steps solve new tasks. It has been particularly successful in computer vision classification and reinforcement learning settings (Nichol et al., 2018). Within NLP, MAML has been adapted to a wide spectrum of supervised problems—including text classification, natural language inference, question answering, summarisation and named entity recognition—where a pre-trained encoder such as BERT is further fine-tuned on small datasets (Rajeswaran et al., 2019; Raghu et al., 2021; Hou et al., 2022). These studies therefore operate (i) on encoder-only, masked-language models and (ii) at parameter counts close to the original 110M-parameter BERT. They leave open whether optimisation-based meta-learning helps decoder LMs and whether its benefits persist at larger parameter scales.
Meta-learning for pretraining. Initial NLP attempts applied MAML only at fine-tuning scale (Raghu et al., 2021; Hou et al., 2022). More recent work embeds bilevel objectives directly in pre-training (Miranda et al., 2023; Ke et al., 2021). While promising, these efforts evaluate only a single model size, focus on one downstream task, or release neither code nor weights, limiting reproducibility and obscuring scale trends. We embed meta-learning directly into the pretraining loop, evaluate on various unseen domains in an unseen task, and provide open weights (11M-570M) and layer-wise spectra, filling that gap.
Subset-Mask LMs (SMLMT). SMLMT constructs pseudo-tasks using a subset of vocabulary words (Bansal et al., 2020). Given an unlabeled text corpus, one selects a set of words and builds an -way classification task. For each chosen word, sentences containing it are collected and the word is masked out. The task is then to predict the masked word from the candidates. Li and Zhang (2021) interleaves it with ProtoNet tasks; we interleave with vanilla LM updates and scale to 570M params.
Interpretable training dynamics. Various works discuss the training of language models in phase transitions Olsson et al. (2022); Hoogland et al. (2024), describing broad changes in indicators as the model gains rapidly in capabilities over a short period of time. We study such phase transitions in the context of meta-learning in pretraining.
@@ 3.4 Downstream protocol
Named entity recognition (NER), the downstream task for this study, is a fundamental NLP task that identifies and categorizes entities (e.g., persons, organizations, locations) within unstructured text (Chinchor and Robinson, 1997), and is used in healthcare (Kundeti et al., 2016; Polignano et al., 2021; Shafqat et al., 2022), law (Leitner et al., 2019; Au et al., 2022; Naik et al., 2023), business (Putthividhya and Hu, 2011; Alvarado et al., 2015; Zhao et al., 2021), and knowledge graph systems Al-Moslmi et al. (2020). Specifically, we evaluate our models on Universal NER benchmark (Mayhew et al., 2024). UNER v1 comprises three categories of NER evaluation data, each built on top of Universal Dependencies (UD) (Nivre et al., 2016, 2020) tokenization and annotations: publicly available in-language treebanks, parallel UD (PUD) evaluation, and other eval-only sets (Appendix B).
After pretraining we load the checkpoint at step 6000 and attach a fresh linear classifier for UniversalNER. Two fine-tuning settings are used: head-only and full. In the head-only setting the Transformer is frozen so fine-tuning mirrors the inner loop, in the full setting all weights update. Fine-tuning uses AdamW at for at most ten epochs with early stopping on development F1.
@@ 4 Model Pretraining
The prerequisite for modifying a pretraining method is ensuring the model still learns. All four Pico-MAML variants reach their respective vanilla loss 1.3–1.6× sooner; perplexity improves only in the 11 M model, indicating an optimisation–regularisation trade-off.
@@
| pico-decoder-tiny-1 | 5.31 | 786.85 |
| pico-maml-decoder-tiny-1 | 4.44 | 422.42 |
| pico-decoder-small-1 | 4.14 | 80.25 |
| pico-maml-decoder-small-1 | 3.67 | 113.76 |
| pico-decoder-medium-1 | 3.89 | 77.90 |
| pico-maml-decoder-medium-2 | 3.49 | 78.63 |
| pico-decoder-large-1 | 3.69 | 49.86 |
| pico-maml-decoder-large-1 | 3.49 | 66.62 |
Contrary to expectation, MAML’s inductive bias may favor optimization over regularization. MAML accelerates convergence but degrades out-of-task fluency at medium+ scales.
@@ 5 Downstream NER Evaluation
Models are fine-tuned on each dataset in Universal NER (Mayhew et al., 2024; Nivre et al., 2016, 2020) with publicly available train and dev sets 333Namely, ddt, ewt, set, bosque, snk, set, talbanken, gsd, gsdsimp, all. Results (averaged across each finetuning dataset) are shown as micro-F1 scores in Table 3, organized by evaluation group: seen (language with full train/test/dev splits), test-only (using Parallel Universal Dependencies PUD), and test-only low-resource languages (e.g., Cebuano, Tagalog). Heatmaps of individual models are available in Appendix C.3.
@@
The most striking takeaway from this stage is that, when averaged across all evaluation steps in a category, absolute F1 remains low () due to poor zero-shot transfer, especially for logographic scripts. Overall, MAML improves mean F1 by 2-3 points at medium/large scales, confirming a modest “learning-to-learn” effect under full adaptation.444While these results are much worse in comparison to the baseline in the original Universal NER paper (Mayhew et al., 2024), this is likely because is a multilingual model (Conneau et al., 2020) and the pretraining dataset for Pico is entirely in English..
@@
To better understand how meta-initialization influences cross-lingual transfer on seen languages, F1 scores are broken down by dataset within the in-language group. The results are separated by tuning regime to clarify the extent to which meta-learned representations help when only the classifier is updated (head-only) versus when the entire model is fine-tuned.
In the head-only setting (Table 7), absolute F1 scores remain low across most datasets. Tiny models fail to generalize altogether. As seen in MAML shows the strongest and most consistent gains at large scales (Table 4)—most prominently on en_ewt, hr_set, and sv_talbanken-suggesting that episodic pretraining creates more adaptable feature spaces, particularly for common entity types and scripts. On Chinese (zh_gsd, zh_gsdsimp), performance is uniformly poor, confirming the baseline result in Mayhew et al. (2024) that transfer from phonographic to logographic scripts is difficult.
@@
In the full setting (Table 5), both vanilla and MAML-pretrained models achieve higher F1 scores across the board. MAML confers consistent +0.01-0.03 gains at medium and large scales, especially for structurally complex languages like Croatian. These relative gains grow as model capacity increases, indicating that larger models benefit more from MAML pretraining. Even in Chinese, where scores are lowest, MAML nudges performance upward. These gains confirm that meta-pretraining does more than support shallow transfer: it reshapes the optimization landscape of the full model in a way that accelerates convergence and improves generalization.
Taken together, these tables validate that MAML pretraining injects a scalable and tunable learning-to-learn signal. However, these average metrics do not tell the full story. Some settings, entity classes, and fine-tuning conditions benefit substantially more than others.
@@
Meta-pretraining yields a clear capacity threshold in head-only adaptation. Under a frozen backbone, only the large model consistently converts its learned initialization into PER (+0.034) and LOC (+0.023) gains; medium and smaller variants lack the representational bandwidth to rewire person and place distinctions via a shallow classifier. By contrast, even medium and small models see gains in ORG (+0.016 F1) likely because organization names often include distinctive tokens (e.g., “Inc.”, “Corp.”, or “University”) that form rigid, token-level co-occurrence patterns. These simple patterns mirror the pseudo-classification episodes SMLMT generates, so a shallow classifier can latch onto them without requiring deep feature reconfiguration.
Full fine-tuning broadens and amplifies these effects. In the full setting, PER sees the largest MAML-induced lift (up to +0.027 in the large model). LOC improvements (+0.016 at large scale) climb more gradually: place names often span heterogeneous contexts and scripts (e.g. Zagreb vs. Beijing), so meta-pretraining must be supplemented by full gradient flow for location-specific embeddings. ORG continues to enjoy gains (+0.012 at large), reinforcing that organization recognition remains the simplest class to bootstrap from episodic tasks.
@@
Now, we discuss how inductive biases manifest in zero-shot cross-lingual transfer to low-resource languages—namely, Tagalog (tl) and Cebuano (ceb).
Tagalog and Cebuano are the two most widely spoken native languages in the Philippines, with tens of millions of first-language speakers each. Both are typologically Austronesian and low-resource, but differ significantly. Tagalog is a morphologically rich, predicate-initial language with a complex voice system that encodes syntactic roles (agent, patient, locative, etc.) through verbal affixes and aspect-marking (Kroeger, 1993; Schachter and Otanes, 1983; Ramos, 2021). Word order is flexible and often pragmatically driven, which weakens the utility of positional cues for tasks like named entity recognition. Cebuano is similarly Austronesian but morphologically simpler than Tagalog, with fewer voice alternations and less affixal variation (Tanangkingsing, 2011). It also does not consistently mark syntactic roles with overt case particles; entities must be inferred from context rather than surface markers (Sityar, 2000). Additionally, Cebuano exhibits a distinct orthographic tradition and more conservative vocabulary (e.g., less Spanish borrowing) (Bunye and Yap, 1971), which further distances it from the English-centric token distributions that dominate cross-lingual pretraining datasets. These characteristics make them ideal stress tests for testing the inductive bias of pretraining strategies like MAML.
@@
In the head-only setting, MAML delivers its greatest impact on small and medium models. For example, the small head jumps from 0.088 to 0.221 overall—an absolute gain of 0.133 F1—and sees particularly large lifts in Cebuano (+0.153) and Tagalog-TRG (+0.262). The medium head also benefits substantially, improving from 0.259 to 0.322. Even the large head picks up a modest +0.030 F1. Only the tiny head collapses, reflecting its inability to form reliable prototypes during meta-training. These patterns suggest that MAML’s episodic learning instills useful, language-agnostic representations in the classifier layers, enabling mid-size heads to generalize token-level cues to new languages without modifying the backbone.
Once we allow full fine-tuning, however, most of MAML’s advantages disappear at higher capacities. The small model retains a small +0.026 F1 edge, but the medium shows no net change and the large actually drops by 0.034. This reversal implies that when every parameter is free to update, the strong gradient signals of full fine-tuning quickly override the meta-learned inductive biases, erasing or even inverting MAML’s earlier head-only gains. The tiny model again underperforms, consistent with its tendency to overfit during meta-training when unconstrained by a fixed backbone.
In the UNER benchmark, Tagalog and Cebuano serve as canonical low-resource, typologically distinct evaluation settings. Overall NER performance remains modest, but, as Table 6 shows, MAML provides meaningful zero-shot boosts in the head-only regime for small and medium models. These gains suggest that even without training exposure to these languages, the inductive biases from English episodic training transfer surprisingly well, at least for token-level prototypes.

### v2

@@ Abstract
Large language models are powerful but costly. We ask whether meta-learning can make the pretraining of small language models not only faster but also more interpretable. We integrate first–order MAML with subset-masked LM pretraining, producing four LLama-style decoder-only models (11M–570M params), and evaluate on multilingual Universal NER. Compared with vanilla training, our hybrid setup (i) reaches the same loss up to 1.6× sooner, (ii) yields modest but consistent average gains on Universal NER at medium/large scales under equal compute (+2–3 percentage points), and (iii) and (iii) reveals phase-like learning dynamics: models first diversify their representations, then compress them in a pattern that aligns with improved episodic accuracy. These observations are correlational, not causal, and we do not claim generality beyond NER or across seeds. We also document a trade-off: perplexity on Paloma (a diverse language modeling benchmark spanning 18 domains; Magnusson et al. (2024)) is worse at most scales. Code, checkpoints and analysis logs are released.
davidafrica/pico-maml
DavidDemitriAfrica/pico-maml-train
@@
David Demitri Africa††thanks: Corresponding author: david.demitri.africa@gmail.com Yuval Weiss Paula Buttery Richard Diehl Martinez University of Cambridge
@@ 1 Introduction
Small language models (SLMs) are attractive for privacy and energyreasons, but trail large models partly because they converge slowly and plateau early (Godey et al., 2024; Biderman et al., 2023; Diehl Martinez et al., 2024). As opposed to the common method of brute-force scaling, we explore a different axis: learning rules. First-order Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) promises a learn-to-learn initialization, yet has rarely been applied to decoder models, and its effect on learning dynamics are poorly understood.We address this by adding meta-learning in model pretraining111Using a lightweight modification of pico-train (Diehl Martinez, 2025), a language model pretraining suite., interleaving ordinary next-token loss (keeps fluency) with 32-way subset-mask (Bansal et al., 2020; Li and Zhang, 2021) episodes (forces rapid binding). Only a tiny MLP head isadapted in the inner loop, so we can track backbone weightswithout gradient noise. Our contributions are:
@@
Four open SLMs (11M → 570M) trained with this hybrid MAML rule.
@@
A public trainer that logs per-checkpoint singular-value spectra, head entropies and query accuracy.
@@
A candid evaluation on Universal NER: modest gains at medium/large scales (+2–3 pp), alongside a perplexity trade-off.
4.
Observational evidence of a diversify–then–compress phase transition in effective rank.
Reporting and scope.
All pretraining and fine-tuning results are from a single shared seed per condition due to compute limits; we therefore report averages across datasets where applicable, avoid statistical claims, and treat learning-dynamics findings as exploratory. We limit generalization claims to NER and to our training regime.
@@ 2 Related Work
Meta-learning for NLP.(MAML; Finn et al., 2017) is an optimisation‑based form of meta‑learning that learns an initialisation from which a few gradient steps solve new tasks. It has been particularly successful in computer vision classification and reinforcement learning settings (Nichol et al., 2018). Within NLP, MAML has been adapted to a wide spectrum of supervised problems—including text classification, natural language inference, question answering, summarisation and named entity recognition—where a pre‑trained encoder such as BERT is further fine‑tuned on small datasets (Rajeswaran et al., 2019; Raghu et al., 2021; Hou et al., 2022). These studies operate (i) on encoder‑only, masked‑language models and (ii) at parameter counts close to the original 110M-parameter BERT. They leave open whether optimisation‑based meta-learning helps decoder LMs and whether its benefits persist at larger parameter scales.Meta-learning for pretraining.Initial NLP attempts applied MAML only at fine-tuning scale(Raghu et al., 2021; Hou et al., 2022). More recent work embedsbilevel objectives directly in pre-training(Miranda et al., 2023; Ke et al., 2021). While promising, these efforts evaluate only a single model size, focus on one downstream task, or release neither code nor weights, limiting reproducibility and obscuring scale trends. We embed meta-learning directly into the pretraining loop, evaluate on various unseen domains in an unseen task, and provide open weights (11M-570M) and layer-wise spectra, filling that gap.Subset-Mask LMs (SMLMT).SMLMT constructs pseudo-tasks using a subset of vocabulary words (Bansal et al., 2020). Given an unlabeled text corpus, one selects a set of words and builds an -way classification task. For each chosen word, sentences containing it are collected and the word is masked out. The task is then to predict the masked word from the candidates. Li and Zhang (2021) interleaves it with ProtoNet tasks; we interleave with vanilla LM updates and scale to 570M params.Interpretable training dynamics.Various works discuss the training of language models in phase transitions Olsson et al. (2022); Hoogland et al. (2024), describing broad changes in indicators as the model gains rapidly in capabilities over a short period of time. We study such phase transitions in the context of meta-learning in pretraining.Effective-rank probes (entropy of singular values) highlight learning behavior in deep nets (Diehl Martinez et al., 2024). Lower-rank structure and rank compression are well documented in the literature (Huh et al., 2021; Galanti et al., 2022; Jaderberg et al., 2014), and we focus on the timing and co-evolution of the measurements of effective-rank probes with episodic generalization under the hybrid objective (§5).
@@ 3.4 Downstream protocol
Named entity recognition (NER), the downstream task for this study, is a fundamental NLP task that identifies and categorizes entities (e.g., persons, organizations, locations) within unstructured text (Chinchor and Robinson, 1997), and is used in healthcare (Kundeti et al., 2016; Polignano et al., 2021; Shafqat et al., 2022), law (Leitner et al., 2019; Au et al., 2022; Naik et al., 2023), business (Putthividhya and Hu, 2011; Alvarado et al., 2015; Zhao et al., 2021), and knowledge graph systems Al-Moslmi et al. (2020). Specifically, we evaluate our models on Universal NER benchmark (Mayhew et al., 2024). UNER v1 comprises three categories of NER evaluation data, each built on top of Universal Dependencies (UD) (Nivre et al., 2016, 2020) tokenization and annotations: publicly available in-language treebanks, parallel UD (PUD) evaluation, and other eval-only sets (Appendix B).After pretraining we load the checkpoint at step 6000 and attach a fresh linear classifier for UniversalNER. Two fine-tuning settings are used: head-only and full. In the head-only setting the Transformer is frozen so fine-tuning mirrors the inner loop, in the full setting all weights update. Fine-tuning uses AdamW at for at most ten epochs with early stopping on development F1.
@@ 4 Model Pretraining
The prerequisite for modifying a pretraining method is ensuring the model still learns. All four Pico-MAML variants reach their respective vanilla loss 1.3–1.6× sooner (faster optimization), but Paloma perplexity is worse at most scales by 6000 steps (Table 2).
@@
| pico-decoder-tiny | 5.31 | 786.85 |
| pico-maml-decoder-tiny | 4.44 | 422.42 |
| pico-decoder-small | 4.14 | 80.25 |
| pico-maml-decoder-small | 3.67 | 113.76 |
| pico-decoder-medium | 3.89 | 77.90 |
| pico-maml-decoder-medium | 3.49 | 78.63 |
| pico-decoder-large | 3.69 | 49.86 |
| pico-maml-decoder-large | 3.49 | 66.62 |
Contrary to expectation, MAML’s inductive bias may favor optimization over regularization. MAML accelerates convergence but degrades out-of-task fluency at most scales. However, this pattern is consistent with known multi-task interference: the episodic discriminative objective improves adaptation signals but can conflict with next-token distributional modeling under fixed compute and a single set of hyperparameters Kendall et al. (2017); Yu et al. (2020); Standley et al. (2020). Hence, it is unclear if the perplexity gap is an objective-mixing artifact or evidence that meta-learning inherently harms LM fluency.
@@ 5 Downstream NER Evaluation
Models are fine-tuned on each dataset in Universal NER (Mayhew et al., 2024; Nivre et al., 2016, 2020) with publicly available train and dev sets 333Namely, ddt, ewt, set, bosque, snk, set, talbanken, gsd, gsdsimp, all. Results (averaged across each finetuning dataset) are shown as micro-F1 scores in Table 3, organized by evaluation group: seen (language with full train/test/dev splits), test-only (using Parallel Universal Dependencies PUD), and test-only low-resource languages (e.g., Cebuano, Tagalog). We report delta F1 as percentage points (pp) unless explicitly marked as percent change (%).
@@
The most striking takeaway from this stage is that, when averaged across all evaluation steps in a category, absolute F1 remains low (, i.e., ) due to poor zero-shot transfer, especially for logographic scripts. Overall, MAML improves mean uplift is approximately +2–3 pp when averaged over all in-language datasets at medium/large scales, confirming a modest “learning-to-learn” effect under full adaptation444While these results are much worse in comparison to the baseline in the original Universal NER paper (Mayhew et al., 2024), this is likely because is a multilingual model (Conneau et al., 2020) and the pretraining dataset for Pico is entirely in English..
@@
To better understand how meta-initialization influences cross-lingual transfer on seen languages, F1 scores are broken down by dataset within the in-language group. The results are separated by tuning regime to clarify the extent to which meta-learned representations help when only the classifier is updated (head-only) versus when the entire model is fine-tuned.In the head-only setting (Table 7), absolute F1 scores remain low across most datasets. Tiny models fail to generalize altogether. MAML shows the strongest and most consistent gains at large scales (Table 4)—most prominently on en_ewt, hr_set, and sv_talbanken-suggesting that episodic pretraining creates more adaptable feature spaces, particularly for common entity types and scripts. On Chinese (zh_gsd, zh_gsdsimp), performance is uniformly poor, confirming the baseline result in Mayhew et al. (2024) that transfer from phonographic to logographic scripts is difficult.
@@
In the full setting (Table 5), both vanilla and MAML-pretrained models achieve higher F1 scores across the board. MAML confers consistent +0.01-0.03 gains at medium and large scales, especially for structurally complex languages like Croatian. These relative gains grow as model capacity increases, indicating that larger models benefit more from MAML pretraining. Even in Chinese, where scores are lowest, MAML nudges performance upward. These gains confirm that meta-pretraining does more than support shallow transfer: it reshapes the optimization landscape of the full model in a way that accelerates convergence and improves generalization.Taken together, these tables validate that MAML pretraining injects a scalable and tunable learning-to-learn signal. However, these average metrics do not tell the full story. Some settings, entity classes, and fine-tuning conditions benefit substantially more than others.
@@
Meta‐pretraining yields a clear capacity threshold in head‐only adaptation. Under a frozen backbone, only the large model consistently converts its learned initialization into PER (+0.034) and LOC (+0.023) gains; medium and smaller variants lack the representational bandwidth to rewire person and place distinctions via a shallow classifier. By contrast, even medium and small models see gains in ORG (+0.016 F1) likely because organization names often include distinctive tokens (e.g., “Inc.”, “Corp.”, or “University”) that form rigid, token‐level co‐occurrence patterns. These simple patterns mirror the pseudo-classification episodes SMLMT generates, so a shallow classifier can latch onto them without requiring deep feature reconfiguration.Full fine‐tuning broadens and amplifies these effects. In the full setting, PER sees the largest MAML‐induced lift (up to +0.027 in the large model). LOC improvements (+0.016 at large scale) climb more gradually: place names often span heterogeneous contexts and scripts (e.g. Zagreb vs. Beijing), so meta‐pretraining must be supplemented by full gradient flow for location‐specific embeddings. ORG continues to enjoy gains (+0.012 at large), reinforcing that organization recognition remains the simplest class to bootstrap from episodic tasks.
@@
Now, we discuss how inductive biases manifest in zero‐shot cross‐lingual transfer to low‐resource languages—namely, Tagalog (tl) and Cebuano (ceb).Tagalog and Cebuano are the two most widely spoken native languages in the Philippines, with tens of millions of first-language speakers each. Both are typologically Austronesian and low-resource, but differ significantly. Tagalog is a morphologically rich, predicate-initial language with a complex voice system that encodes syntactic roles (agent, patient, locative, etc.) through verbal affixes and aspect-marking (Kroeger, 1993; Schachter and Otanes, 1983; Ramos, 2021). Word order is flexible and often pragmatically driven, which weakens the utility of positional cues for tasks like named entity recognition. Cebuano is similarly Austronesian but morphologically simpler than Tagalog, with fewer voice alternations and less affixal variation (Tanangkingsing, 2011). It also does not consistently mark syntactic roles with overt case particles; entities must be inferred from context rather than surface markers (Sityar, 2000). Additionally, Cebuano exhibits a distinct orthographic tradition and more conservative vocabulary (e.g., less Spanish borrowing) (Bunye and Yap, 1971), which further distances it from the English-centric token distributions that dominate cross-lingual pretraining datasets. These characteristics make them ideal stress tests for testing the inductive bias of pretraining strategies like MAML.
@@
In the head-only setting, MAML delivers its greatest impact on small and medium models. For example, the small head jumps from 0.088 to 0.221 overall—an absolute gain of 0.133 F1—and sees particularly large lifts in Cebuano (+0.153) and Tagalog-TRG (+0.262). The medium head also benefits substantially, improving from 0.259 to 0.322. Even the large head picks up a modest +0.030 F1. Only the tiny head collapses, reflecting its inability to form reliable prototypes during meta-training. These patterns suggest that MAML’s episodic learning instills useful, language-agnostic representations in the classifier layers, enabling mid-size heads to generalize token-level cues to new languages without modifying the backbone.Once we allow full fine-tuning, however, most of MAML’s advantages disappear at higher capacities. The small model retains a small +0.026 F1 edge, but the medium shows no net change and the large actually drops by 0.034. This reversal implies that when every parameter is free to update, the strong gradient signals of full fine-tuning quickly override the meta-learned inductive biases, erasing or even inverting MAML’s earlier head-only gains. The tiny model again underperforms, consistent with its tendency to overfit during meta-training when unconstrained by a fixed backbone.In the UNER benchmark, Tagalog and Cebuano serve as canonical low-resource, typologically distinct evaluation settings. Overall NER performance remains modest, but, as Table 6 shows, MAML provides meaningful zero-shot boosts in the head-only regime for small and medium models. These gains suggest that even without training exposure to these languages, the inductive biases from English episodic training transfer surprisingly well, at least for token-level prototypes.